A solution S of sulfuric acid is prepared by dissolving 49g of H2SO4 in 500 mL of distilled water. Calculate the molar concentration of the solution


1
Expert's answer
2021-01-11T03:51:12-0500

The density of H2SO4 is 1.83 g/cm3. Therefore, the volume of H2SO4 equals "\frac{49g}{1.83g/cm^3}=27mL" , and the total volume of the resulting solution is 500 + 27 = 527 mL = 0.527 L.

Amount of H2SO4 in moles:


"n=\frac{49g}{98.1g/mol}=0.50 mol"


Therefore, the concentration is:


"C=\frac{0.50mol}{0.527L}=0.95M"


Answer: 0.95 M